OpenStudy (anonymous):

Given f′′(x)=−9sin(3x) and f′(0)=−4 and f(0)=0. Find f(π/5)=

OpenStudy (anonymous):

I got f'(x)=-3cos(3x)+c c=-1 f(x)=sin(3x)-x-D D=0 F(pi/5)=0.3227

OpenStudy (anonymous):

the correct answer is -3.447

OpenStudy (anonymous):

I don't know what i did wrong

OpenStudy (zehanz):

If f''(x)=-9sin(3x) then f'(x)=acos(3x)+c Now differentiate acos(3x) to get a: -3asin(3x)=-9sin(3x), so a=3.

OpenStudy (anonymous):

what is acos?

OpenStudy (zehanz):

acos is a*cos

OpenStudy (nurali):

f''(x)=−9sin(3x) f'(x) = 3cos(3x)+C -4=3+C C=-7 f'(x) = 3cos(3x)-7 f(x) = sin(3x)-7x+D D=0 f(x) = sin(3x)-7x+0 f(π/5)=sin(3*π/5)-7*π/5 f(π/5)=_______

OpenStudy (zehanz):

So your f' is right. Now f'(x)=3cos(3x)+c, f'(0)=-4, so 3cos(0)+c=-4, so 3+c=-4, so c=-7. f(x)=sin(3x) -7x + d. f(0)=0: sin(0)-7*0+d=0, so d=0 f(x)=sin(3x)-7x. f(pi/5)=sin(3pi/5)-7pi/5 which is about 0.951- 4.398=-3.447

OpenStudy (zehanz):

You had -x instead of -7x in f.

OpenStudy (anonymous):

oh ok